tywanikia:

Given 649-04-04-00-00_files/i0160000.jpg find 649-04-04-00-00_files/i0160001.jpg. a. 649-04-04-00-00_files/i0160002.jpg c. 649-04-04-00-00_files/i0160003.jpg b. 649-04-04-00-00_files/i0160004.jpg d. 649-04-04-00-00_files/i0160005.jpg

Vocaloid:

from what I can tell it's asking for the inverse of g(x) = 3/(x^2 + 2x) change g(x) to y to get y = 3/(x^2 + 2x) then switch all the x's with y's and vice versa to get x = 3/(y^2+2y) then solve again for y